An experiment shows that a 260 mL gas sample has a mass of 0.429 g at a pressure of 765 mm Hg and a temperature of 22 C. what is the molar mass of the gas?
Answer #1

volume of gas= 260 ml= 0.26 L
pressure = 765 mm Hg=1.00658 atm
temperature T = 22 C+273 = 295 K
mass of gas m= 0.429 g
R= 0.0821 Latm/mol/k
Molar mass M =??
use ideal gas equation
PV =nRT

PV = m*RT/M

thus molar mass M = m*RT/PV = 0.429 g*0.0821 Latm/mol/k*295 K/1.00658 atm*0.26 L
= 39.7 g/mol
********************
:)